the Kb of ammonia (NH3) at 25.0 C is 1.75 x 10-5. Calculate the pH of a 0.0321 M aqueous solution of NH4Cl
Answer #1

NH4Cl(aq) -------------> NH4^+ (aq) + Cl^- (aq)

0.0321M                        0.0321M

             NH4^+ (aq) + H2O(l) --------------> NH3(aq) + H3O^+ (aq)

I            0.0321                                              0                  0

C           -x                                                     +x               +x

E        0.0321-x                                               +x              +x

                Ka   = Kw/Kb

                       = 1*10^-14/(1.75*10^-5)

                        = 5.7*10^-10

                 Ka    = [NH3][H3O^+]/[NH4^+]

               5.7*10^-10 = x*x/(0.0321-x)

               5.7*10^-10*(0.0321-x) = x^2

                 x   = 4.27*10^-6

              [H3O^+]   = x   = 4.27*10^-6M

                PH   = -log[H3O^+]

                          = -log4.27*10^-6

                           = 5.3696>>>>>answer